OpenStudy (anonymous):

Add and simplify: x^-1 + x^-2

OpenStudy (anonymous):

x^-1 =1/x x^-2=1/x^2

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

x^-1 + x^-2 =1/x + 1/x^2

OpenStudy (anonymous):

now, simplify it

OpenStudy (anonymous):

got it